OpenStudy (anonymous):

Two boat landings are 1.0 km apart on the same bank of a stream that flows at 1.4 km/h. A motorboat makes the round trip between the two landings in 50 minutes. What is the speed of the boat relative to the water?

OpenStudy (abhisar):

Here again you have to use the same concept.

OpenStudy (abhisar):



OpenStudy (abhisar):

Let the speed of the boat wrt the river be x, then actual speed of boat would become (x+1.4) and (x-1.4).

OpenStudy (abhisar):

Can you find the value of x now ?

OpenStudy (anonymous):

I'm not sure I understand this one

OpenStudy (abhisar):

you mean you don't understand ?

OpenStudy (anonymous):

yeah

OpenStudy (abhisar):

Did u understood upto what i wrote above ?

OpenStudy (anonymous):

yes, i understand what you drew, but I'm still confused on how to find the value of x

OpenStudy (abhisar):

brb !

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar Now, total time taken to cover the round trip is 50 mins = 0.83 hours. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

Time = Distance/speed. So, \(\huge\sf \frac{1}{(x+1.4)}+\frac{1}{(x-1.4)}=0.83\) Solve it and find the value of x

OpenStudy (abhisar):

Getting it ?

OpenStudy (anonymous):

yeah, so far

OpenStudy (abhisar):

can you calculate now ?

OpenStudy (abhisar):

Remember a^2 - b^2 = (a - b)(a + b)

OpenStudy (anonymous):

so would I have to multiply both sides by (x+1.4)(x-1.4) to get rid of the fractions

OpenStudy (anonymous):

Then I would get \[(x+1.4)+(x-1.4)=.83(x+1.4)(x-1.4)\]?

OpenStudy (abhisar):

\(\sf (x+1.4)(x-1.4)=x^2-(1.4)^2\)

OpenStudy (abhisar):

I used the identity \(\sf a^2-b^2=(a+b)(a-b)\)

OpenStudy (anonymous):

but I'm still solving for x, right?

OpenStudy (abhisar):

yes. Add the left hand side of the equation@Abhisar \(\huge\sf \frac{1}{(x+1.4)}+\frac{1}{(x-1.4)}=0.83\)

OpenStudy (anonymous):

so the answer I get when I do that problem is \[x ^{2}-1.96\]

OpenStudy (abhisar):

It will be the denominator of the left hand side after adding. Find numerator too

OpenStudy (abhisar):

\(\huge\sf \frac{(x+1.4)+(x-1.4)}{x ^{2}-1.96}=0.83\) Now find x

OpenStudy (anonymous):

wouldn't it be easier to factor the bottom then find x, or can you not do that?

OpenStudy (abhisar):

Do however you like. I find this method easier. The answer to your question will be the value of x

OpenStudy (anonymous):

okay so I would multiply both sides by \[x ^{2}-1.96\] to get \[(x ^{2}-1.96)+(x+1.4)+(x-1.4)=0.83(x ^{2}-1.96)\] right?

OpenStudy (anonymous):

Then I would add everything up on the left hand side and get \[(x ^{2}+2x-1.96)=.83x ^{2}-1.6268\]

OpenStudy (anonymous):

That's where I get stuck

OpenStudy (abhisar):

\(\huge\sf \frac{(x+1.4)+(x-1.4)}{x ^{2}-1.96}=0.83\) --> 2x= \(\sf 0.83x^2-1.62\) -->\(\sf 0.83x^2-2x-1.62=0\) Finally you will have to solve the quadratic equation to get the value of x. `This is the easiest way.`

OpenStudy (anonymous):

okay, hold on

OpenStudy (anonymous):

okay I set it up as \[x=\frac{ 2\pm \sqrt{(-2)^{2}-4(0.83)(-1.62)} }{ 2(.83) }\] teh when I type it all in, my answer is \[x=\frac{ 2\pm \sqrt{9.3784} }{ 1.66 }\]

OpenStudy (abhisar):

it will be 3 m/s. But do cross check.

OpenStudy (anonymous):

yeah, thats what I got

OpenStudy (anonymous):

So, that's it? The speed is 3m/s?

OpenStudy (abhisar):

yes i think so.

OpenStudy (anonymous):

what if I had the same speed boats but the distance was 6 instead of 1. How would that affect the equation?

OpenStudy (abhisar):

Do the same way

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar yes. Add the left hand side of the equation@Abhisar \(\huge\sf \frac{1}{(x+1.4)}+\frac{1}{(x-1.4)}=0.83\) \(\color{blue}{\text{End of Quote}}\) just put 6 at the place of 1 in this equation.

OpenStudy (anonymous):

and it also says the measurement is in km/h, so it's actually not 3m/s it's still 3km/h?

OpenStudy (anonymous):

it says that's the right answer. thank you so much!

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)